OpenStudy (eem0102):

Let f(x) = x + 8 and g(x) = x2 − 6x − 7. Find f(g(2)).

OpenStudy (mathmale):

This is a "composite function" problem. Determine the value of g(2). Use this value as the input to function f(x). That'll give you the answer you want.

OpenStudy (mathmale):

@eem0102: As a matter of etiquette: If you post a problem, please stick with it for a while instead of logging out. Others may want to help you solve your math problem yourself, but not to do the work for you.

OpenStudy (michael_t):

f(x)=-7

OpenStudy (mathmale):

Let f(x) = x + 8 and g(x) = x2 − 6x − 7. Find f(g(2)). First, evaluate the function g(x) = x^2 - 6x - 7 at x=2. Please show your work.